If the sides of a square are increased by 3 ​inches, the area becomes 81 square inches. Find the length of the sides of the original square.
Found 2 solutions by Alan3354, Boreal:
Answer by Alan3354(69443)

If the sides of a square are increased by 3 ​inches, the area becomes 81 square inches. Find the length of the sides of the original square.
-------------
Area = 81
Find the length of the sides.
subtract 3.

Answer by Boreal(15235)

side is x
x^2 is area
make side x+3
area is (x+3^2)=x^2+6x+9=81
x^2+6x-72=0
(x+12)(x-6)=0
only positive root is x=6 inches,
sides of original were 6 x 6 inches
when they become 9 x 9, the area is 81 sq inches.